# Fourier Learning Machines: Nonharmonic Fourier-Based Neural Networks for Scientific Machine Learning

### v1

@@ Abstract
We introduce the Fourier Learning Machine (FLM), a neural network (NN) architecture designed to represent a multidimensional nonharmonic Fourier series. The FLM uses a simple feedforward structure with cosine activation functions to learn the frequencies, amplitudes, and phase shifts of the series as trainable parameters. This design allows the model to create a problem–specific spectral basis adaptable to both periodic and nonperiodic functions. Unlike previous Fourier–inspired NN models, the FLM is the first architecture able to represent a complete, separable Fourier basis in multiple dimensions using a standard Multilayer Perceptron–like architecture. A one–to–one correspondence between the Fourier coefficients and amplitudes and phase-shifts is demonstrated, allowing for the translation between a full, separable basis form and the cosine phase–shifted one. Additionally, we evaluate the performance of FLMs on several scientific computing problems, including benchmark Partial Differential Equations (PDEs) and a family of Optimal Control Problems (OCPs). Computational experiments show that the performance of FLMs is comparable, and often superior, to that of established architectures like SIREN and vanilla feedforward NNs.
Keywords: Nonharmonic Fourier Series, Multilayer Perceptron, Partial Differential Equations, Evolutionary Games, Rock–Paper–Scissors.
@@ 1 Introduction
Motivated by Joseph Fourier’s early $19^{th}$ century idea that any arbitrary function, regardless of its complexity or continuity, could be represented as a sum of sinessines and cosinescosines, and by the successful application of NNs as universal function approximators since the late $20^{th}$ century, we propose in this paper a new Fourier–inspired NN model, and we shall name it the Fourier Learning Machine (FLM). Unlike typical black–box NN models, the FLM’s internal structure is built to directly represent an $m$-dimensional nonharmonic Fourier series. It achieves this by relying solely on the foundational components of a classical multilayer perceptron (MLP): namely, affine transformations followed by nonlinear activation functions (in the context of this paper, these being pure cosinescosines). This design choice eschews more complicated architectural features, such as multiplication nodes, recurrent relations or any other customized components, previously found in predecessor models. With hardwired integer weights in its input–to–hidden layer, the FLM represents an $m$-dimensional, full-basis classical (finite) Fourier series as a special case. In order to tackle the approximation of nonperiodic functions, we relax the integrality constraints of the frequency components, thus configuring the frequencies as learnable parameters of the FLM. This allows for the search for the sinusoidal basis functions that best represent the target map. While maintaining interpretability, FLMs build a function–specific spectral representation for a wide variety of functions, including nonperiodic ones. As such, our proposed model combines the interpretability of Fourier series with the data–driven nature of NNs (and, in particular, the simplicity of the MLP architecture), effectively representing a nonharmonic Fourier series [1, 2] in which the basis functions are learned and adapted to the target function it is meant to approximate.
@@
    We provide a comprehensive review of all previous attempts to construct such a function approximator, typically referred in the literature as “Fourier Neural Networks”, departing from foundational works [3] and covering up to recent architectures with trigonometric activation functions [4]. Within this context, we demonstrate that the FLM is not only architecturally simpler than previous models but also computationally competitive, often outperforming established, traditional NN models.
@@ 2 Related Literature
The idea of a Fourier series–inspired NN was first proposed in the late 1980s, just before NNs were shown to possess universal function approximation properties [5]. This pioneer architecture, called “Fourier Network” (FN) was proposed by Gallant and White in 1988 [3]. In their work, they carefully constructed a Fourier series representation into a NN by defining a “cosinecosine–squasher” activation function, a domain–constrained, monotonic, cosinecosine function. By fixing the input–to–hidden layer weights, this NN was demonstrated to mimic a finite multidimensional Fourier series. This “cosinecosine–squasher” is a monotonic and piecewise differentiable function, $F:\mathbb{R}\rightarrow[0,1]$ , defined as
@@
While $F(\cdot)$ is not strictly periodic, the FN constructs the Fourier cosinecosine bases by constraining and summing multiple hidden units within a group of neurons called “clumps”. This activation function is designed to fit the “squashing” paradigm of early NNs [6] while allowing for the construction of a NN replicating the approximation properties of Fourier series. A single “clump” is depicted in Fig. 1. As such, the output of each of these “clumps” forms one of the cosinecosine or sinesine basis (by adding a phase–shift) present in a Fourier series.
Figure 1: A “clump” of neurons that outputs the $i^{th}$ single cosinecosine basis of the Fourier Network appearing in Gallant and White [3].
To create a single cosinecosine basis, the weighted sum of the inputs, i.e., the pre–activation $net_{i}=\boldsymbol{w}_{i}^{\top}\boldsymbol{x}+b_{i}$ is preprocessed and activated using the cosinecosine–squasher function. By constraining $\xi_{im}=\delta_{im}=2\theta_{i}$, where $i=1,2,\dots,K$ and $m=-M_{i},\dots,0,\dots,M_{i}$, the output of the FN is given by,
@@
which is a classical Fourier series in cosinecosine phase-shifted form when an extra term $\delta_{0}=\theta_{0}+\sum_{i=1}^{K}\left[\theta_{i}-\theta_{i}\,2(2M_{i}+1)\right]$ is added to the output. The role of $M_{i}$ is to bound the range within which the output $net_{i}$ can vary. As the input $\boldsymbol{x}$ varies in the (perturbed) hypercube $\epsilon,2\pi-\epsilon]^{m}$, $net_{i}$ varies in the interval $2\pi M_{i},2\pi(M_{i}+1)]$. Therefore, $M_{i}$ is inherently dependent on the “hardwired” parameters $\mathbf{w}_{i}$ and $b_{i}$ in the first layer. Furthermore, $M_{i}$ increases with $i$, that is, the inclusion of higher frequency modes requires more neurons within a “clump”. For smaller network sizes, while the authors claim that by undoing the “hardwiring” (in $\mathbf{w}_{i}$ and $b_{i}$) and removing the constraint in $\xi_{im}$ and $\delta_{im}$ the FN could potentially perform better, doing so nullifies its exact correspondence with a finite Fourier series: making $\mathbf{w}_{i}$ and $b_{i}$ trainable implies that one could not determine a specific value for each $M_{i}$ before training the FN. Thus, it is unclear how to pick $M_{i}$ under the aforementioned construction. Lastly, by construction, removing the constraint $\xi_{im}=\delta_{im}=2\theta_{i}$ would hinder the $i^{th}$ “clump” to create a cosinecosine basis. Therefore, such an attempt would not lead the network to represent a nonharmonic Fourier series.
@@
It is worth noticing that the requirement of a “monotonic” activation function was later dropped based on the universal approximation (and convergence) result in [7], which established that nonpolynomial and bounded functions (e.g., cosinecosine) can be also serve as activation functions in NNs. Therefore, such a convoluted design proposed in [3] can be partially justified by the lack of such a result when their architecture was developed.
@@
In the context of systems control and identification, [8] proposed the first NN capable of mimicking a finite Fourier series by using pure sinessines and cosinescosines, naming it the “Fourier Series Neural Network” (FSNN). Despite their NN capability to represent a full-basis Fourier series, given a fixed set of nontrainable “frequency weights”, the architecture chosen relies on multiplicative nodes to compose each of the basis functions. The training process in this case is rather simple, given the linearity (with respect to the Fourier coefficients) of the output layer, thus yielding a linear optimization problem. More recently, [9] introduced an adaptive controller where a FSNN is used to compute and adjust the gains of a PID controller online.
@@
With an emphasis on similarities of its model with the brain neurophysiology, [10] proposed an architecture relying on multiplicative nodes that attempted to reproduce a finite Fourier series in cosinecosine phase–shifted form (optionally “squeezed” through a sigmoid function). This network, named “Fourier Neural Network” (the first appearance of such a term in the literature) assumed the form
@@
However, it is implicit in [10] that the trainable parameters $w$ and $\varphi$ are not fixed or constrained to be either natural numbers (including zero) or odd–multiple of $\frac{\pi}{2}$, respectively. Therefore, despite its similarities with a Fourier series, $f(\cdot)$ fails to represent either a finite, classical Fourier series or a (typical) nonharmonic one. In its current form, a classical Fourier series could be attained from (2) by fixing $w$’s to a predetermined set of integers and making some $\varphi\in\{(2k+1)\frac{\pi}{2}\;|\;k\in\mathbb{Z}\}$. Conversely, a nonharmonic series can only be obtained in the form of a sum of cosinecosine–only basis functions, which would configure a very specific (and arguably weak) form of nonharmonic Fourier series, allowing one to approximate even functions only. For a depiction of this FN model, the reader is pointed to [11].
@@
In [12], a single–input, multiple–output NN is proposed, wherein each output corresponds to that of an individual finite Fourier series. In this architecture, all the series share the same basis functions but each have their own set of (trainable) Fourier coefficients. Besides being of limited applicability (once it is design to approximate single variable functions only), [12] also fails to make any reference to previously proposed, Fourier–inspired NN already in existence in the literature at that time.
@@
The authors of [13] use a Discrete Fourier Transform to find the hidden–to–output weights for the FSNN model from [8]. This neural network formulation, however, is unnecessary for the task, as the calculation provides no advantage over a direct computation performed outside of an NN framework.
@@
In [14], another type of FNN is proposed, this version utilizing a vector form of the multidimensional, finite Fourier series. This NN assumes the form
$f(\boldsymbol{x})=\sum_{i=1}^{N}\Bigg{[}a_{i}\,\sin\Big{(}\boldsymbol{c}_{i}^{\top}\boldsymbol{x}\Big{)}+b_{i}\,\cos\Big{(}\boldsymbol{d}_{i}^{\top}\boldsymbol{x}\Big{)}\Bigg{]},$ (3)
@@
where, the (real–valued) numbers $\boldsymbol{c}_{i}$, $\boldsymbol{d}_{i}$, $a_{i}$ and $b_{i}$ are the learnable parameters of the corresponding NN. Unless the frequency components are constrained to be equal position-wise, i.e., $\boldsymbol{c}_{i}=\boldsymbol{d}_{i}$, this representation is a trigonometric expansion, but it is not strictly a nonharmonic Fourier series. This is because the use of different frequency vectors inside the sinesine and cosinecosine prevents (3) to be written as its equivalent exponential or separable basis forms. Therefore, in its original form presented, (3) represents neither a classical Fourier series nor a nonharmonic one.
@@
A comparative study on the performance of the FN model of [3], the FNN of [10] and the FNN of [14] is presented in [11], which concluded that the top performing model among these three is the early FN model in [3]. However, this FN model of [3] was shown to underperform a vanilla feedforward NN with sigmoid activation function. The reason behind this result, as hypothesized by the authors, is that, outside the interval $\frac{\pi}{2},\frac{\pi}{2}]$, the “cosinecosine–squasher” as written in (1) is constant, whereas the sigmoid employed in the vanilla feedforward NN is asymptotically constant outside of this same interval.
@@
An investigation on the use of sinesine functions as activation functions in NNs first appeared in [15]. In this work, the authors claim that the wave–like behavior introduced in the loss function by the presence of sinessines introduces many local minima, making such models harder to train. However, the output of the NNs chosen for the experiments shown in [15] are not meant to represent a Fourier series in any form. Furthermore, the performance of such activation function is not studied for a simple feedforward NN structure, but in the context of encoder–decoder architectures based on recurrent and long short–term memory NNs. Related to this idea, [16] designs a “Sinusoidal Neural Network (SNN)”, with a newly defined sinesine–based node operations inspired by information propagation in the brain. Even though empirical results presented are promising, no theoretical background is provided to guarantee the convergence of this NN.
@@
A (deep) NN with sinesine activation functions, called SIREN (Sinusoidal Representation Network) is introduced in [4], tailored for applications in digital media processing. The authors customize the first layer of this NN in order to extract high–frequency components of the target data (image, video or audio), doing so by multiplying the weighted sum of the input by a relatively large constant term $\omega_{0}$ (keeping the bias untouched), while all the other activation functions assume the form $y\longmapsto\sin{y}$. We note that a SIREN–1, that is, a SIREN with one single layer, for $\omega_{0}=1$ can be shown, under certain trigonometric identities, to assume the mathematical form of
$f(\mathbf{x})=\sum_{i=1}^{N}\Bigg{[}a_{i}\,\sin\Big{(}\boldsymbol{w}_{i}^{\top}\boldsymbol{x}\Big{)}+b_{i}\,\cos\Big{(}\boldsymbol{w}_{i}^{\top}\boldsymbol{x}\Big{)}\Bigg{]}.$ (4)
@@
Furthermore, we note that (4) is identical to (3) for when $\mathbf{c}_{i}=\mathbf{d}_{i}$. Thus, similar to the case of [14], a SIREN–1 represents a Fourier series by fixing the components of $\mathbf{w}$ as integers (noted only later on by [17]), and a nonharmonic series otherwise. It is worth noting that in [4] there is no mention to any potential relations between a SIREN NN and any form of a Fourier series. However, as it is also the case of the FNN of [14], this Fourier series cannot be represented in a full–basis, separable form.
@@
In the context of scientific machine learning, [18] introduced a one–dimensional FNN as function approximators. Their FNN model assumes the form of a nonharmonic (phase–shifted cosinecosine) Fourier series (with trainable frequency components). Nonetheless, all the numerical experiments reported considered only the approximation of periodic functions that can be fully decomposed as a finite Fourier series, thus not fully exploring the potential of their model in the approximation of nonperiodic functions.
@@
In order to avoid confusion by the use of similar terms and concepts, other scientific contributions that somehow involve the use of Fourier analysis (but do not attempt to construct a NN that replicates a Fourier series) must be mentioned in order to avoid confusion. Recently, the Fourier Neural Operator (FNO) gained tremendous attention in the scientific machine learning community [19]. The FNO primary goal is to approximate a operator, that is, a mapping between function spaces. Thus, it is suitable to learn operators in the context of approximating the solution of entire parametric families of PDEs. In this regard, the “Fourier” comes from the fact that the integral kernel is parameterized in Fourier space. The FLM model presented in this paper, instead, aims to approximate a mapping from a higher–dimension Euclidean space to a real number. In [20], preprocessing input vectors with a Fourier feature mapping allows this NN to overcome its inherent spectral bias, i.e., the tendency to learn low-frequency components and therefore facilitating learning of high-frequency functions. Finally, by automatically determining the optimal sampling frequency and number of neurons, Fourier-based extreme learning machines are applied in [21] for the solution of PDEs.
@@
Some of the models described above move toward learnable spectral bases, while others are constrained by the integrality of their frequency components. Nonetheless, in either case, they often lack a systematic architecture allowing them to represent a complete, separable, and interpretable Fourier series of the target function (either classical or nonharmonic). Even though SIREN–1 and the FNN in [18] fulfill the aforementioned criteria for a single input, these properties don’t hold true when considering the general, multidimensional case. As we will show later in this paper, FLMs fill this gap, being the first Fourier–replicating NN that is capable of representing both finite and nonharmonic Fourier series in a full, separable basis form in $m$ dimensions by using only MLP–like, feedforward NN operations.

### v3

# Fourier Learning Machines: Nonharmonic Fourier–Based Neural Networks for Scientific Machine Learning

@@ Abstract
We introduce the Fourier Learning Machine (FLM), a neural network (NN) architecture designed to represent a multidimensional nonharmonic Fourier series. The FLM uses a simple feedforward structure with cosine activation functions to learn the frequencies, amplitudes, and phase shifts of the series as trainable parameters. This design allows the model to create a problem–specific spectral basis adaptable to both periodic and nonperiodic functions. Unlike previous Fourier–inspired NN models, the FLM is the first architecture able to represent a multidimensional Fourier series with a complete set of basis functions in separable form, doing so by using a standard Multilayer Perceptron–like architecture. A one–to–one correspondence between the Fourier coefficients and amplitudes and phase-shifts is demonstrated, allowing for the translation between a full, separable basis form and the cosine phase–shifted one. Additionally, we evaluate the performance of FLMs on several scientific computing problems, including benchmark Partial Differential Equations (PDEs) and a family of Optimal Control Problems (OCPs). Computational experiments show that the performance of FLMs is comparable, and often superior, to that of established architectures like SIREN and vanilla feedforward NNs.
@@ 1 Introduction
Motivated by Joseph Fourier’s early $19^{th}$ century idea that any arbitrary function, regardless of its complexity or continuity, could be represented as a sum of sinessines and cosinescosines, and by the successful application of NNs as universal function approximators since the late $20^{th}$ century, we propose in this paper a new Fourier–inspired NN model, and we shall name it the Fourier Learning Machine (FLM). Unlike typical black–box NN models, the FLM’s internal structure is built to directly represent an $m$-dimensional nonharmonic Fourier series. It achieves this by relying solely on the foundational components of a classical multilayer perceptron (MLP): namely, affine transformations followed by nonlinear activation functions (in the context of this paper, these being pure cosinescosines). This design choice eschews more complicated architectural features, such as multiplication nodes, recurrent relations or any other customized components, previously found in predecessor models. With hardwired integer weights in its input–to–hidden layer, the FLM represents an $m$-dimensional, full-basis classical (finite) Fourier series as a special case. In order to tackle the approximation of nonperiodic functions, we relax the integrality constraints of the frequency components, thus configuring the frequencies as learnable parameters of the FLM. This allows for the search for the sinusoidal basis functions that best represent the target map. While maintaining interpretability, FLMs build a function–specific spectral representation for a wide variety of functions, including nonperiodic ones. As such, our proposed model combines the interpretability of Fourier series with the data–driven nature of NNs (and, in particular, the simplicity of the MLP architecture), effectively representing a nonharmonic Fourier series (Young, 1981; Duffin and Schaeffer, 1952) in which the basis functions are learned and adapted to the target function it is meant to approximate.
@@
    We provide a comprehensive review of all previous attempts to construct such a function approximator, typically referred in the literature as “Fourier Neural Networks”, departing from foundational works (Gallant and White, 1988) and covering up to recent architectures with trigonometric activation functions (Sitzmann et al., 2020). Within this context, we demonstrate that the FLM is not only architecturally simpler than previous models but also computationally competitive, often outperforming established, traditional NN models.
@@ 2 Related Literature
The idea of a Fourier series–inspired NN was first proposed in the late 1980s, just before NNs were shown to possess universal function approximation properties (Hornik et al., 1989). This pioneer architecture, called “Fourier Network” (FN) was proposed by Gallant and White (1988). In their work, they carefully constructed a Fourier series representation into a NN by defining a “cosinecosine–squasher” activation function, a domain–constrained, monotonic, cosinecosine function. By fixing the input–to–hidden layer weights, this NN was demonstrated to mimic a finite multidimensional Fourier series. This “cosinecosine–squasher” is a monotonic and piecewise differentiable function, $F:\mathbb{R}\rightarrow[0,1]$ , defined as
@@
While $F(\cdot)$ is not strictly periodic, the FN constructs the Fourier cosinecosine bases by constraining and summing multiple hidden units within a group of neurons called “clumps”. This activation function is designed to fit the “squashing” paradigm of early NNs (Rumelhart et al., 1986) while allowing for the construction of a NN replicating the approximation properties of Fourier series. A single “clump” is depicted in Fig. 1. As such, the output of each of these “clumps” forms one of the cosinecosine or sinesine basis (by adding a phase–shift) present in a Fourier series.
Figure 1: A “clump” of neurons that outputs the $i^{th}$ single cosinecosine basis of the Fourier Network appearing in Gallant and White (1988).
To create a single cosinecosine basis, the weighted sum of the inputs, i.e., the pre–activation $net_{i}=\bm{w}_{i}^{\top}\bm{x}+b_{i}$ is preprocessed and activated using the cosinecosine–squasher function. By constraining $\xi_{im}=\delta_{im}=2\theta_{i}$, where $i=1,2,\dots,K$ and $m=-M_{i},\dots,0,\dots,M_{i}$, the output of the FN is given by,
@@
which is a classical Fourier series in cosinecosine phase-shifted form when an extra term $\delta_{0}=\theta_{0}+\sum_{i=1}^{K}\left[\theta_{i}-\theta_{i}\,2(2M_{i}+1)\right]$ is added to the output. The role of $M_{i}$ is to bound the range within which the output $net_{i}$ can vary. As the input $\bm{x}$ varies in the (perturbed) hypercube $\epsilon,2\pi-\epsilon]^{m}$, $net_{i}$ varies in the interval $2\pi M_{i},2\pi(M_{i}+1)]$. Therefore, $M_{i}$ is inherently dependent on the “hardwired” parameters $\mathbf{w}_{i}$ and $b_{i}$ in the first layer. Furthermore, $M_{i}$ increases with $i$, that is, the inclusion of higher frequency modes requires more neurons within a “clump”. For smaller network sizes, while the authors claim that by undoing the “hardwiring” (in $\mathbf{w}_{i}$ and $b_{i}$) and removing the constraint in $\xi_{im}$ and $\delta_{im}$ the FN could potentially perform better, doing so nullifies its exact correspondence with a finite Fourier series: making $\mathbf{w}_{i}$ and $b_{i}$ trainable implies that one could not determine a specific value for each $M_{i}$ before training the FN. Thus, it is unclear how to pick $M_{i}$ under the aforementioned construction. Lastly, by construction, removing the constraint $\xi_{im}=\delta_{im}=2\theta_{i}$ would hinder the $i^{th}$ “clump” to create a cosinecosine basis. Therefore, such an attempt would not lead the network to represent a nonharmonic Fourier series.
@@
It is worth noticing that the requirement of a “monotonic” activation function was later dropped based on the universal approximation (and convergence) result in Leshno et al. (1993), which established that nonpolynomial and bounded functions (e.g., cosinecosine) can be also serve as activation functions in NNs. Therefore, such an intricate design proposed in Gallant and White (1988) can be partially justified by the lack of such a result when their architecture was developed.
@@
In the context of systems control and identification, Zhu and Paul (1995) proposed the first NN capable of mimicking a finite Fourier series by using pure sinessines and cosinescosines, naming it the “Fourier Series Neural Network” (FSNN). Despite their NN capability to represent a full-basis Fourier series, given a fixed set of nontrainable “frequency weights”, the architecture chosen relies on multiplicative nodes to compose each of the basis functions. The training process in this case is rather simple, given the linearity (with respect to the Fourier coefficients) of the output layer, thus yielding a linear optimization problem. More recently, Benrabah et al. (2021) introduced an adaptive controller where a FSNN is used to compute and adjust the gains of a PID controller online.
@@
With an emphasis on similarities of its model with the brain neurophysiology, Silvescu (1999) proposed an architecture relying on multiplicative nodes that attempted to reproduce a finite Fourier series in cosinecosine phase–shifted form (optionally “squeezed” through a sigmoid function). This network, named “Fourier Neural Network” (the first appearance of such a term in the literature) assumed the form
@@
However, it is implicit in Silvescu (1999) that the trainable parameters $w$ and $\varphi$ are not fixed or constrained to be either natural numbers (including zero) or odd–multiple of $\frac{\pi}{2}$, respectively. Therefore, despite its similarities with a Fourier series, $f(\cdot)$ fails to represent either a finite, classical Fourier series or a (typical) nonharmonic one. In its current form, a classical Fourier series could be attained from Equation 2 by fixing $w$’s to a predetermined set of integers and making some $\varphi\in\{(2k+1)\frac{\pi}{2}\;|\;k\in\mathbb{Z}\}$. Conversely, a nonharmonic series can only be obtained in the form of a sum of cosinecosine–only basis functions, which would configure a very specific (and arguably weak) form of nonharmonic Fourier series, allowing one to approximate even functions only. For a depiction of this FN model, the reader is pointed to Uteuliyeva et al. (2020).
@@
In Yang et al. (2003), a single–input, multiple–output NN is proposed, wherein each output corresponds to that of an individual finite Fourier series. In this architecture, all the series share the same basis functions but each have their own set of (trainable) Fourier coefficients. Besides being of limited applicability (once it is design to approximate single variable functions only), Yang et al. (2003) also fails to make any reference to previously proposed, Fourier–inspired NN already in existence in the literature at that time.
@@
Halawa (2008) uses a Discrete Fourier Transform to find the hidden–to–output weights for the FSNN model from Zhu and Paul (1995). This neural network formulation, however, is unnecessary for the task, as the calculation provides little advantage over a direct computation performed outside of an NN framework.
@@
In Liu (2013), another type of FNN is proposed, this version utilizing a vector form of the multidimensional, finite Fourier series. This NN assumes the form
$f(\bm{x})=\sum_{i=1}^{N}\Bigg[a_{i}\,\sin\Big(\bm{c}_{i}^{\top}\bm{x}\Big)+b_{i}\,\cos\Big(\bm{d}_{i}^{\top}\bm{x}\Big)\Bigg],$ (3)
@@
where, the (real–valued) numbers $\bm{c}_{i}$, $\bm{d}_{i}$, $a_{i}$ and $b_{i}$ are the learnable parameters of the corresponding NN. Unless the frequency components are constrained to be equal position-wise, i.e., $\bm{c}_{i}=\bm{d}_{i}$, this representation is a trigonometric expansion, but it is not strictly a nonharmonic Fourier series. This is because the use of different frequency vectors inside the sinesine and cosinecosine prevents Equation 3 to be written as its equivalent exponential or separable basis forms. Therefore, in its original form presented, Equation 3 represents neither a classical Fourier series nor a nonharmonic one.
@@
A comparative study on the performance of the FN model of Gallant and White (1988), the FNN of Silvescu (1999) and the FNN of Liu (2013) is presented in Uteuliyeva et al. (2020), which concluded that the top performing model among these three is the early FN model of Gallant and White (1988). However, this FN model of Gallant and White (1988) was shown to underperform a vanilla feedforward NN with sigmoid activation function. The reason behind this result, as hypothesized by the authors, is that, outside the interval $\frac{\pi}{2},\frac{\pi}{2}]$, the “cosinecosine–squasher” as written in Equation 1 is constant, whereas the sigmoid employed in the vanilla feedforward NN is asymptotically constant outside of this same interval.
@@
A comprehensive and critical investigation on the use of sinesine functions as activation functions in NNs has appeared in Parascandolo et al. (2017). In this work, the authors claim that the wave–like behavior introduced in the loss function by the presence of sinessines introduces many local minima, making such models harder to train. However, the output of the NNs chosen for the experiments shown in Parascandolo et al. (2017) are not meant to represent a Fourier series in any form. Furthermore, the performance of such activation function is not studied for a simple feedforward NN structure, but in the context of encoder–decoder architectures based on recurrent and long short–term memory NNs. Related to this idea, Ozmermer (2020) designs a “Sinusoidal Neural Network (SNN)”, with a newly defined sinesine–based node operations inspired by information propagation in the brain. Even though empirical results presented are promising, no theoretical background is provided to guarantee the convergence of this NN.
@@
A (deep) NN with sinesine activation functions, called SIREN (Sinusoidal Representation Network) is introduced in Sitzmann et al. (2020), tailored for applications in digital media processing. The authors customize the first layer of this NN in order to extract high–frequency components of the target data (image, video or audio), doing so by multiplying the weighted sum of the input by a relatively large constant term $\omega_{0}$ (keeping the bias untouched), while all the other activation functions assume the form $y\longmapsto\sin{y}$. We note that a SIREN–1, that is, a SIREN with one single layer, for $\omega_{0}=1$ can be shown, under certain trigonometric identities, to assume the mathematical form of
$f(\mathbf{x})=\sum_{i=1}^{N}\Bigg[a_{i}\,\sin\Big(\bm{w}_{i}^{\top}\bm{x}\Big)+b_{i}\,\cos\Big(\bm{w}_{i}^{\top}\bm{x}\Big)\Bigg].$ (4)
@@
Furthermore, we note that Equation 4 is identical to Equation 3 for when $\mathbf{c}_{i}=\mathbf{d}_{i}$. Thus, similar to the case of Liu (2013), a SIREN–1 represents a Fourier series by fixing the components of $\mathbf{w}$ as integers (noted only later on by Benbarka et al. (2022)), and a nonharmonic series otherwise. It is worth noting that in Sitzmann et al. (2020) there is no mention to any potential relations between a SIREN NN and any form of a Fourier series. However, as it is also the case of the FNN of Liu (2013), this Fourier series cannot be represented in a full–basis, separable form.
@@
In the context of scientific machine learning, Ngom and Marin (2021) introduced a one–dimensional FNN as function approximators. Their FNN model assumes the form of a nonharmonic (phase–shifted cosinecosine) Fourier series (with trainable frequency components). Nonetheless, all the numerical experiments reported considered only the approximation of periodic functions that can be fully decomposed as a finite Fourier series, thus not fully exploring the potential of their model in the approximation of nonperiodic functions.
@@
In order to avoid confusion by the use of similar terms and concepts, other scientific contributions that somehow involve the use of Fourier analysis (but do not attempt to construct a NN that replicates a Fourier series) must be mentioned in order to avoid confusion. Recently, the Fourier Neural Operator (FNO) gained tremendous attention in the scientific machine learning community (Li et al., 2020). The FNO primary goal is to approximate a operator, that is, a mapping between function spaces. Thus, it is suitable to learn operators in the context of approximating the solution of entire parametric families of PDEs. In this regard, the “Fourier” comes from the fact that the integral kernel is parameterized in Fourier space. The FLM model presented in this paper, instead, aims to approximate a mapping from a higher–dimension Euclidean space to a real number. In Tancik et al. (2020), preprocessing input vectors with a Fourier feature mapping allows this NN to overcome its inherent spectral bias, i.e., the tendency to learn low–frequency components and therefore facilitating learning of high–frequency functions. Lastly, by automatically determining the optimal sampling frequency and number of neurons, Fourier–based extreme learning machines are applied in Ma et al. (2022) for the solution of PDEs.
@@
Some of the models described above move toward learnable spectral bases, while others are constrained by the integrality of their frequency components. Nonetheless, in either case, they often lack a systematic architecture allowing them to represent a complete, separable, and interpretable Fourier series of the target function (either classical or nonharmonic). Even though SIREN–1 and the FNN in Ngom and Marin (2021) fulfill the aforementioned criteria for a single input, these properties don’t hold true when considering the general, multidimensional case. As we will show later in this paper, FLMs fill this gap, being the first Fourier–replicating NN that is capable of representing both finite and nonharmonic Fourier series in a full, separable basis form in $m$ dimensions by using only MLP–like, feedforward NN operations.
Lastly, even though explicitly outside of the NN domain (at least in the sense in which NNs are used in this work, i.e., for deterministic function approximation), it is worth mentioning that FLMs share a fundamental similarity with Sparse Spectral Gaussian Processes (SSGP), introduced in Lázaro-Gredilla et al. (2010). This relationship draws from established results bridging NNs and Gaussian Processes (GPs) (Neal (1996); Murphy (2023)). Both FLMs and SSGP rely on learning frequencies of their underlying trigonometric basis functions. However, FLMs rely on the usual gradient descent in the space of frequencies (and amplitudes) to find the set of frequencies that “best” describe the target function. As it will be shown in Section 5, this is not necessarily a data-driven approach, but instead based on the customized loss function that enforces the desired system dynamics. On the other hand, in SSGP, the frequencies are treated as hyperparameters and optimized by maximizing the marginal likelihood (thus data-driven) to approximate the spectral density of the covariance kernel. Therefore, SSGP falls into the framework of (efficient) probabilistic regression for large-scale, discrete data, not undermining the FLM’s spectral adaptability, but rather reinforcing its importance and applicability.